# V-Max: Making RL Practical for Autonomous Driving

### v1

@@ Abstract
Learning-based decision-making has the potential to enable generalizable Autonomous Driving (AD) policies, reducing the engineering overhead of rule-based approaches. Imitation Learning (IL) remains the dominant paradigm, benefiting from large-scale human demonstration datasets, but it suffers from inherent limitations such as distribution shift and imitation gaps. Reinforcement Learning (RL) presents a promising alternative, yet its adoption in AD remains limited due to the lack of standardized and efficient research frameworks. To this end, we introduce V-Max, an open research framework providing all the necessary tools to make RL practical for AD. V-Max is built on Waymax Gulino et al. (2023), a hardware-accelerated AD simulator designed for large-scale experimentation. We extend it using ScenarioNet’s Li et al. (2023b) approach, enabling the fast simulation of diverse AD datasets. V-Max integrates a set of observation and reward functions, transformer-based encoders, and training pipelines. Additionally, it includes adversarial evaluation settings and an extensive set of evaluation metrics. Through a large-scale benchmark, we analyze how network architectures, observation functions, training data, and reward shaping impact RL performance.
@@ 1 Introduction
However, applying RL to real-world tasks such as AD introduces significant challenges, particularly regarding sample efficiency and training environments. As a result, RL remains underused in AD research due to practical constraints. Imitation Learning (IL, Bansal et al. (2019)) is often favored instead, as it capitalizes on vast driving datasets collected by vehicle fleets and reduces decision-making to a supervised learning task. The absence of RL-compatible environments made RL unusable in the only public challenge for AD Karnchanachari et al. (2024), which led the organizers to conclude that learning-based methods could not compete with simple rule-based baselines Dauner et al. (2023).
@@ 3 The V-Max Framework
Figure 1 provides an overview of the V-Max framework, which formulates mid-to-end AD as a partially observable Markov Decision Process (POMDP, Spaan (2012)). In this section, we present the core components of V-Max and how they extend Waymax to make RL practical for AD.
@@ 5.1 Standard evaluation
Table 6: V-Max benchmark, a complete definition of metrics is provided in Appendix A.
@@
Collision rate $\downarrow$ 0.56 7.50 4.70 1.74 2.16 7.81 13.14
Offroad $\downarrow$ 0.75 3.79 1.44 0.74 1.04 1.14 6.92
Red-light violation $\downarrow$ 0.61 0.54 0.49 0.09 0.19 0.33 0.55
Progress $\uparrow$ 97.30 151.00 158.00 169.01 159.61 189.52 86.87
At-fault collisions $\downarrow$ 0.49 3.25 1.11 1.17 1.46 5.57 4.49
Making progress $\uparrow$ 99.41 96.90 96.87 98.37 97.87 97.58 87.97
Driving direction compliance $\uparrow$ 97.91 99.30 99.37 98.75 97.90 99.26 97.47
TTC within bound $\uparrow$ 99.21 91.99 94.58 92.48 95.24 83.63 94.79
Progress along route ratio $\uparrow$ 99.09 92.76 92.68 95.49 93.32 94.61 78.64
Speed limit compliance $\uparrow$ 99.97 85.38 85.89 99.53 99.78 98.35 98.55
Multiple lanes score $\uparrow$ 91.26 99.08 96.91 93.00 93.00 94.00 89.68
Comfort $\uparrow$ 80.16 53.74 33.58 53.18 52.74 18.69 93.32
Accuracy $\uparrow$ 98.06 88.20 93.40 97.44 96.61 90.75 79.42
V-Max score $\uparrow$ 0.93 0.81 0.82 0.88 0.86 0.78 0.72

### v2

# V-Max: A Reinforcement Learning Framework for Autonomous Driving

@@ Abstract
Learning-based decision-making has the potential to enable generalizable Autonomous Driving (AD) policies, reducing the engineering overhead of rule-based approaches. Imitation Learning (IL) remains the dominant paradigm, benefiting from large-scale human demonstration datasets, but it suffers from inherent limitations such as distribution shift and imitation gaps. Reinforcement Learning (RL) presents a promising alternative, yet its adoption in AD remains limited due to the lack of standardized and efficient research frameworks. To this end, we introduce V-Max, an open research framework providing all the necessary tools to facilitate RL research for AD. V-Max is built on Waymax Gulino et al. (2023), a hardware-accelerated AD simulator designed for large-scale experimentation. We extend it using ScenarioNet’s Li et al. (2023b) approach, enabling the fast simulation of diverse AD datasets. V-Max integrates a set of observation and reward functions, transformer-based encoders, and training pipelines. Additionally, it includes adversarial evaluation settings and an extensive set of evaluation metrics. Through a large-scale benchmark, we investigate how network architectures, observation functions, training data, and reward shaping impact RL performance.
@@ 1 Introduction
However, applying RL to real-world tasks such as AD introduces significant challenges, particularly regarding sample efficiency and training environments. As a result, RL remains underused in AD research due to practical constraints. Imitation Learning (IL, Bansal et al. (2019)) is often favored instead, as it capitalizes on vast driving datasets collected by vehicle fleets and reduces decision-making to a supervised learning task. The absence of RL-compatible environments made RL unusable in the nuPlan challenge Karnchanachari et al. (2024), which led the organizers to conclude that learning-based methods could not compete with simple rule-based baselines Dauner et al. (2023).
@@ 3 The V-Max Framework
Figure 1 provides an overview of the V-Max framework, which formulates mid-to-end AD as a partially observable Markov Decision Process (POMDP, Spaan (2012)). In this section, we present the core components of V-Max and how they extend Waymax to facilitate RL’s application to AD.
@@ 5.1 Standard evaluation
Table 6: V-Max benchmark, evaluations on WOMD Valid, in the closed-loop non-reactive setting. A complete definition of metrics is provided in Appendix A.
@@
Collision rate $\downarrow$ 0.56 7.88 5.95 1.74 2.16 7.81 13.14
Offroad $\downarrow$ 0.75 3.37 1.07 0.74 1.04 1.14 6.92
Red-light violation $\downarrow$ 0.61 0.24 0.21 0.09 0.19 0.33 0.55
Progress $\uparrow$ 97.30 140.44 138.12 169.01 159.61 189.52 86.87
At-fault collisions $\downarrow$ 0.49 3.10 1.00 1.17 1.46 5.57 4.49
Making progress $\uparrow$ 99.41 96.90 96.74 98.37 97.87 97.58 87.97
Driving direction compliance $\uparrow$ 97.91 99.46 99.55 98.75 97.90 99.26 97.47
TTC within bound $\uparrow$ 99.21 93.02 95.99 92.48 95.24 83.63 94.79
Progress along route ratio $\uparrow$ 99.09 91.99 90.84 95.49 93.32 94.61 78.64
Speed limit compliance $\uparrow$ 99.97 100 100 99.53 99.78 98.35 98.55
Multiple lanes score $\uparrow$ 91.26 99.08 96.79 93.00 93.00 94.00 89.68
Comfort $\uparrow$ 80.16 55.25 33.58 53.18 52.74 18.69 93.32
Accuracy $\uparrow$ 98.06 88.54 92.77 97.44 96.61 90.75 79.42
Accuracy only at-fault $\uparrow$ 98.14 93.32 97.72 98.00 97.32 88.07 92.99
V-Max score $\uparrow$ 0.93 0.85 0.85 0.88 0.86 0.78 0.72
@@ 6 Conclusion
Acknowledgments
Use unnumbered third level headings for the acknowledgments. All acknowledgments, including those to funding agencies, go at the end of the paper. Only add this information once your submission is accepted and deanonymized. The acknowledgments do not count towards the 8–12 page limit.
Supplementary Materials
The following content was not necessarily subject to peer review.